OpenStudy (anonymous):

So the question wants me to find out the slope of the line in simplest from for 4x – 8y – 16 = 0 So im GUESSING that two points on the graph would be (4,0) and (0,-12) ?

OpenStudy (anonymous):

form"

OpenStudy (anonymous):

If thats right then I think ive got the rest covered XD

OpenStudy (misty1212):

I!!the slope of \[ax+by=-c\] is always \[m=-\frac{a}{b}\]

OpenStudy (anonymous):

to make it easy for just solve for y

OpenStudy (misty1212):

in your case \(a=4,b=-8\) so it is not hard to compute

OpenStudy (anonymous):

Well i dont really know how to solve fr y in this kind of equation XD

OpenStudy (anonymous):

So its 4/-8?

OpenStudy (misty1212):

do you know how to compute \(-\frac{a}{b}\)?

OpenStudy (misty1212):

oh no dear you are off by a minus sign

OpenStudy (anonymous):

so its 1/-2?

OpenStudy (misty1212):

there is one in the form, and one in the \(-8\)

OpenStudy (anonymous):

didivde top and bottom by bruhh

OpenStudy (anonymous):

i mean

OpenStudy (anonymous):

4

OpenStudy (misty1212):

literally \[-\frac{4}{-8}\] butin reality\[\frac{1}{2}\]

OpenStudy (anonymous):

oh! ooooh ok i get it thanks!

OpenStudy (anonymous):

Thanks both XD

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]